# Large Language Models are Clinical Reasoners: Reasoning-Aware Diagnosis Framework with Prompt-Generated Rationales

### v1

@@ Abstract
Machine reasoning has made great progress in recent years owing to large language models (LLMs). In the clinical domain, however, most NLP-driven projects mainly focus on clinical classification or reading comprehension, and under-explore clinical reasoning for disease diagnosis due to the expensive rationale annotation with clinicians. In this work, we present a “reasoning-aware” diagnosis framework that rationalizes the diagnostic process via prompt-based learning in a time- and labor-efficient manner, and learns to reason over the prompt-generated rationales. Specifically, we address the clinical reasoning for disease diagnosis, where the LLM generates diagnostic rationales providing its insight on presented patient data and the reasoning path towards the diagnosis, namely Clinical Chain-of-Thought (Clinical CoT). We empirically demonstrate LLMs/LMs’ ability of clinical reasoning via extensive experiments and analyses on both rationale generation and disease diagnosis in various settings. We further propose a novel set of criteria for evaluating machine-generated rationales’ potential for real-world clinical settings, facilitating and benefiting future research in this area.
@@ Introduction
To generate clinical CoT rationales, which deliver the diagnostic reasoning towards the correct diagnosis, by prompting a LLM to rationalize the presented clinical data.
@@
*$\mathcal{R}^{*}=\underset{\mathcal{R}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{R}|\mathcal{P},\mathcal{D})$ (4)
@@
Clinical CoT (2-shot) 68.4 77.5 59.3 67.4 76.2 40.5 89.3 62.6 82.2 51.2 60.9 63.8 39.2 87.9
Table 1: Evaluation on LLMs in zero-and-few-shot diagnosis.
@@
Figure 4: Data efficiency brought by clinical reasoning.
@@
(a) Correct Diagnoses
(b) Misdiagnoses
Figure 6: Evaluations on clinical rationales (score range: 0-5).
@@
We observe that GPT-4 and Ref perform similarly within correct diagnoses. However, in Misdiagnoses group, rationales generated without access to ground-truth diagnosis (GPT-4) yield better scores even than those with access (with Ref), in every criterion. This phenomenon indicates that when a ground-truth diagnosis is challenging to predict based on the given patient description (i.e., Misdiagnosis group), it is possible that prior knowledge of the diagnosis may not be beneficial for rationale annotation. We presume it stems from the discrepancy of available clinical evidence between radiologists (who annotated the dataset in the real world) and our study. As a result, when asked to condition on the ground-truth label, GPT-4 may operate as if it is being forced to construct a reasoning path that contradicts its understanding of the available clinical evidence. Therefore, the superior results of rationales generated without referencing to the ground truth in the Misdiagnoses group (GPT-4; yellow bars) manefest that GPT-4 can faithfully reflect the observed clinical evidence in the rationales.
@@
Most importantly, the outstanding human-likeness scores, especially within the correct diagnoses, show that our rationales can effectively replicate the clinical reasoning of radiologists, and thus are more likely to seamlessly integrate into real-world radiology reports.
@@
In conventional data-driven approaches for disease diagnosis, i.e., text or image classification, it is non-trivial to annotate whether a certain feature is important in all possible scenarios. Our reasoning-aware diagnosis framework not only ameliorates this need, but also provides rationales that replicate such mechanism of human radiologist. Hence, we presume our framework has the potential to serve as a reliable tool for assisting clinicians in real-world data annotation.
@@
The success of LMs has sparked a surge in applying NLP techniques to the biomedical field. For example, Lee et al. (2020) fine-tune the commonly used BERT model (Kenton and Toutanova 2019) with medical corpus to endow it with biomedical knowledge, which is then implemented by Yue, Jimenez Gutierrez, and Sun (2020) to solve the clinical reading comprehension task; Rajagopal et al. (2021) and Feng et al. (2023) address the generation of explanations for various medical conditions via sequence-to-sequence language models with template-based approaches. More recently, upon the advancements of LLMs, Agrawal et al. (2022) have proposed to use LLMs to recognize named entities from clinical texts; Li et al. (2023) use LLMs to predict the synergy of drug pairs in rare human tissues that lack structured data and features.
@@ Conclusion
Ethical Consideration and Societal Impacts
@@ Acknowledgements
Appendix A Appendix
Appendix B Details on the Datasets
@@
Appendix C Prompts and Experimental Details

### v3

@@ Abstract
Machine reasoning has made great progress in recent years owing to large language models (LLMs). In the clinical domain, however, most NLP-driven projects mainly focus on clinical classification or reading comprehension, and under-explore clinical reasoning for disease diagnosis due to the expensive rationale annotation with clinicians. In this work, we present a “reasoning-aware” diagnosis framework that rationalizes the diagnostic process via prompt-based learning in a time- and labor-efficient manner, and learns to reason over the prompt-generated rationales. Specifically, we address the clinical reasoning for disease diagnosis, where the LLM generates diagnostic rationales providing its insight on presented patient data and the reasoning path towards the diagnosis, namely Clinical Chain-of-Thought (Clinical CoT). We empirically demonstrate LLMs/LMs’ ability of clinical reasoning via extensive experiments and analyses on both rationale generation and disease diagnosis in various settings. We further propose a novel set of criteria for evaluating machine-generated rationales’ potential for real-world clinical settings, facilitating and benefiting future research in this area. [^1] We release the few-shot exemplars for Clinical Chain-of-Thought at https://github.com/ktio89/ClinicalCoT
@@ Introduction
To generate clinical CoT rationales, which deliver the diagnostic reasoning towards the correct diagnosis, by prompting an LLM to rationalize the presented clinical data.
@@
$\mathcal{R}^{*}=\underset{\mathcal{R}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{R}|\mathcal{P},\mathcal{D})$ (4)
@@
Clinical CoT 68.4 77.5 59.3 67.4 76.2 40.5 89.3 62.6 82.2 51.2 60.9 63.8 39.2 87.9
Table 1: Evaluation on LLMs in zero-and-few-shot diagnosis. The Clinical CoT includes two exemplar shots.
@@
Figure 4: Data efficiency brought by clinical reasoning.
@@
Figure 6: Evaluations on clinical rationales. We report the average score (score range: 0-5).
@@
We observe that GPT-4 and Ref perform similarly within correct diagnoses. However, in Misdiagnoses group, rationales generated without access to ground-truth diagnosis (GPT-4) yield better scores even than those with access to it (i.e., Ref), in every criterion. This phenomenon indicates that when a ground-truth diagnosis is challenging to predict based on the given patient description (i.e., Misdiagnosis group), it is possible that prior knowledge of the diagnosis may not be beneficial for rationale annotation. We presume it stems from the discrepancy of available clinical evidence between radiologists (who annotated the dataset in the real world) and our study. As a result, when asked to condition on the ground-truth label, GPT-4 may operate as if it is being forced to construct a reasoning path that contradicts its understanding of the available clinical evidence. Therefore, the superior results of rationales generated without referencing to the ground truth in the Misdiagnoses group (GPT-4; yellow bars) manefest that GPT-4 can faithfully reflect the observed clinical evidence in the rationales.
@@
Most importantly, the outstanding human-likeness scores, especially within the correct diagnoses, show that our rationales can effectively replicate the clinical reasoning of radiologists, and thus are more likely to be seamlessly integrated into real-world radiology reports.
@@
In conventional data-driven approaches for disease diagnosis, i.e., text or image classification, it is non-trivial to annotate whether a certain feature is important in all possible scenarios. Our reasoning-aware diagnosis framework not only ameliorates this need, but also provides rationales that replicate such mechanism of human radiologists. Hence, we presume our framework has the potential to serve as a reliable tool for assisting clinicians in real-world data annotation.
@@
The success of LMs has sparked a surge in applying NLP techniques to the biomedical field (Lee et al. 2020; Yue, Jimenez Gutierrez, and Sun 2020; Rajagopal et al. 2021; Kim et al. 2023; Feng et al. 2023). For example, Lee et al. (2020) fine-tune the commonly used BERT model (Kenton and Toutanova 2019) with medical corpus to endow it with biomedical knowledge, which is then implemented by Yue, Jimenez Gutierrez, and Sun (2020) to solve the clinical reading comprehension task; Rajagopal et al. (2021) and Feng et al. (2023) address the generation of explanations for various medical conditions via sequence-to-sequence language models with template-based approaches. More recently, upon the advancements of LLMs, Agrawal et al. (2022) have proposed to use LLMs to recognize named entities from clinical texts; Li et al. (2023) use LLMs to predict the synergy of drug pairs in rare human tissues that lack structured data and features.
@@ Conclusion
Ethical Statement
@@ Acknowledgements
Appendix
1 Details on the Datasets
@@
2 Prompts and Experimental Details